# ROSE: Register Assisted General Time Series Forecasting with Decomposed Frequency Learning

### v1

@@ 3.3 Time series register
By decomposed frequency learning, we can obtain the domain-general representations, and additionally, we propose the TS-Register that learns register tokens as the domain-specific information from the multi-domain datasets for adaptive transfer. It clusters domain-specific information from the multi-domain datasets into register tokens and stores such domain-specific information in the register codebook during pre-training. Then, it adaptively selects similar information from the register codebook to enhance the target domain during fine-tuning.
We set up a randomly initialized register codebook $\mathbf{E}\in\mathbb{R}^{H\times D_{\mathrm{r}}}$ with $H$ cluster center vectors $\mathbf{e}_{i}\in\mathbb{R}^{D_{\mathrm{r}}},i\in\{1,2,…,H\}$. Each of input time series $\mathbf{x}\in\mathbb{R}^{L}$ is projected into a data-dependent embedding $\mathbf{x}_{\mathrm{e}}\in\mathbb{R}^{D_{\mathrm{r}}}$ through a linear layer.
Pre-training stage. As shown in Figure 2(b), we use the register codebook to cluster these data-dependent embeddings, which generate domain-specific information, and store them in pre-training. Specifically, We find a cluster center vector $\mathbf{e}_{\delta}$ from the register codebook $\mathbf{E}$ where we use $\delta$ to denote the cluster that the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ belongs to, as shown in Equation 6.
$\delta=\underset{j=1:H}{\arg\operatorname{min}}\left\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{j}\right\|_{2}.$ (6)
$\mathcal{L}_{\text{register}}=\|\text{sg}(\mathbf{x}_{\mathrm{e}})-\mathbf{e}_{\delta}\|^{2}_{2}+\|\mathbf{x}_{\mathrm{e}}-\text{sg}(\mathbf{e}_{\delta})\|^{2}_{2}.$ (7)
To update the cluster center vectors in the register codebook $\mathbf{E}$ that represent the domain information of the pre-trained datasets, we set the loss function shown in Equation 7 that minimizes the distance between the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$. $\text{sg}()$ refers to the stop gradient operation, and is used to solve the problem that the gradient of the $\arg\operatorname{min}$ function cannot be backpropagated [35]. The first term is used to update the register codebook $\mathbf{E}$, and the second term is used to update the parameters of the linear layer that learns $\mathbf{x}_{\mathrm{e}}$.
In this way, the vectors in the register codebook $\mathbf{E}$ cluster the embeddings of different data and learn the domain-specific centers for pre-trained datasets, which can represent domain-specific information. As a vector in the register codebook $\mathbf{E}$, $\mathbf{e}_{\delta}$ represents the domain-specific information for input $\mathbf{x}$. $\mathbf{e}_{\delta}$ is invariant under small perturbations in $\mathbf{x}_{\mathrm{e}}$ that represents $\mathbf{x}$, which promotes better representation of domain-specific information and robustness of the vectors in the register. This also avoids their over-reliance on detailed information about specific datasets [36].
@@
Fine-tuning stage. As shown in Figure 2(c), after obtaining a register codebook $\mathbf{E}$ that contains domain-specific information through pre-training, we freeze the register parameters to adaptively use this domain-specific information in the downstream task.
Since the target domain may not strictly fit one of the upstream domains, learning the embedding of the downstream data employs a top-k strategy that selects top-k similar vectors in the register codebook. As shown in Equation 8, the embedding of input time series $\mathbf{x}_{\mathrm{e}}$ picks the $k$ nearest vectors in the register codebook $\mathbf{E}$, and uses their average as $\bar{\mathbf{e}}_{k}$ to represent the domain-specific information from the pre-train stage. $\bar{\mathbf{e}}_{k}$ is also patched into $\mathbf{X}_{\mathrm{d}}\in\mathbb{R}^{N_{\text{r}}\times D}$ and is used as the register tokens.
@@
Since the downstream data has its own specific information at the dataset level in addition to the domain level, this may not be fully represented by the domain information obtained from the pre-trained dataset alone. Therefore, we additionally set a matrix $\mathbf{A}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ to adjust $\mathbf{X}_{\mathrm{d}}$ to complement the specific information of downstream data. Since the pre-trained model has a very low intrinsic dimension [37], in order to get better fine-tuning results, $\mathbf{A}$ is set as a low-rank matrix:
@@ 3.4 Training
To improve the model’s prediction ability when only few samples are available downstream, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency mask to learn unified features that are more applicable to the downstream prediction task.
[…]
@@
Pre-training. The reconstruction task learns generalized features through the transformer encoder and reconstruction decoder. To utilize these features for the prediction task, the parameters of the reconstruction decoder are copied to the prediction decoder during forward propagation. To avoid prediction training affecting the generalization performance of the model, the gradients of the prediction heads are skipped at back-propagation. The overall pre-training loss is shown in Equation 12.
@@ 4 Experiments
Pre-training datasets. Large-scale datasets are crucial for pre-training a general time series forecasting model. In light of this, we gather a considerable amount of publicly available datasets from various domains, including energy, nature, health, transport, web, and economics, etc. The details of these datasets are shown in the Appendix A.1. To enhance data utilization, we downsampled fine-grained datasets to coarser granularity, resulting in approximately 886 million time points.
@@
Evaluation datasets. To conduct comprehensive and fair comparisons for different models, we conducted experiments on eight real-world benchmarks as the target datasets, including Weather, Solar, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
Baselines. We select the state-of-the-art models as our baselines, including Transformer-based models: Timer [2], iTransformer [20], and PatchTST [6]; CNN-based model: TimesNet [18]; MLP-based models: FITS [38], TIDE [39], and DLinear [40]. It is worth noting that Timer is a recent general time series model with a GPT-style architecture. Since the size of the look-back window can affect the performance of different models, we choose the look-back window size in $\{96,336,512,720\}$ for each baseline that achieves its best results for fair comparisons.
@@
Setup. Consistent with previous works, we adopted Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. All methods predict the future values with lengths $F=\{96,192,336,720\}$. More implementation details are presented in the Appendix A.3.
4.1 Main results
@@
Full-shot setting. We first pre-train ROSE on our pre-training datasets and then fine-tune with full downstream data (full-shot setting). Similarly, all baselines are also run in the same full-shot setting. As shown in Table 1, we also present the results of the ROSE in 10% few-shot setting, where the model uses the same validation and test data but is only fine-tuned with 10% training data of each target dataset. Key observations are summarized as follows. First, as a general forecasting model, ROSE achieves superior performance compared to the seven state-of-the-art baselines with full-data training, especially Timer, which shows that our decomposed frequency learning and register help to learn generalized representations from large-scale datasets and adaptively transfer the multi-domain information to specific downstream scenarios. Second, we observe that ROSE achieves competitive performance in 10% few-shot setting, even compared to baselines trained with full data. This observation validates the transferability of ROSE pre-trained with large multi-source data.
@@
Few-shot setting. We also compared all models under the 10% few-shot setting in Table 9 in Appendix H.2. We can see that ROSE performs the best on all eight datasets in the scenarios where the training data are rare in the target domain, surpassing extensive advanced models. Figure 4 shows the results of pre-trained ROSE and ROSE trained from scratch on the Solar and ETTh1 with different fine-tuning data percentages, noting the best baselines in full-shot setting. The pre-trained ROSE shows stable, superior performance even with limited fine-tuned samples. Specifically, the pre-trained ROSE achieves performance that surpasses the SOTA with only 5% for Solar and 1% data for ETTh1. Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in predictive performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through large-scale pre-training.
Models ROSE ROSE 10% Timer ITransformer PatchTST FITS TimesNet TiDE DLinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 0.341 0.367 0.349 0.372 0.436 0.456 0.407 0.410 0.353 0.382 0.355 0.379 0.400 0.406 0.355 0.378 0.357 0.379
ETTm2 0.246 0.305 0.249 0.308 0.308 0.351 0.288 0.332 0.256 0.317 0.249 0.312 0.291 0.333 0.249 0.312 0.267 0.332
ETTh1 0.391 0.414 0.397 0.419 0.483 0.459 0.454 0.447 0.413 0.434 0.407 0.429 0.458 0.450 0.419 0.430 0.423 0.437
ETTh2 0.330 0.374 0.335 0.380 0.366 0.406 0.383 0.407 0.331 0.381 0.333 0.382 0.414 0.427 0.345 0.394 0.431 0.447
Traffic 0.390 0.264 0.418 0.278 0.413 0.284 0.428 0.282 0.391 0.264 0.410 0.282 0.62 0.336 0.356 0.261 0.434 0.295
Weather 0.217 0.251 0.224 0.252 0.277 0.306 0.246 0.298 0.226 0.264 0.218 0.260 0.259 0.287 0.236 0.282 0.246 0.300
Solar 0.188 0.215 0.192 0.216 0.251 0.262 0.233 0.262 0.207 0.294 0.223 0.264 0.301 0.319 0.223 0.288 0.230 0.295
Electricity 0.155 0.248 0.164 0.253 0.181 0.274 0.178 0.270 0.159 0.253 0.163 0.257 0.192 0.295 0.159 0.257 0.166 0.264
@@ 4.2 Ablation studies
Mask method.
To further validate the effectiveness of decomposed frequency learning, we replace the multi-frequency masking with different masking methods, including two mainstream time-domain methods: patch masking [6] and multi-patch masking [5], as well as a single frequency masking, which masks time series in the frequency domain only once. The results show that frequency domain modeling yields superior performance improvements compared to time-domain modeling. Additionally, we observe that applying multi-frequency masking results in enhanced outcomes, thus affirming that incorporating a multi-frequency perspective significantly aids the model in more accurately comprehending temporal patterns.
Table 2: Ablations on ROSE. We replace different masking methods, in addition to component removal. The average results of all predicted lengths are listed here.

### v2

@@ 3.3 Time series register
By decomposed frequency learning, we can obtain the general representations, and additionally, we propose the TS-Register that learns register tokens as the domain-specific information from the multi-domain datasets for adaptive transfer. It clusters domain-specific information from the multi-domain datasets into register tokens and stores such domain-specific information in the register during pre-training. Then, it adaptively selects domain-specific information from the register via a Top-K selection strategy to enhance the performance in the target domain. A novel learnable low-rank matrix is proposed to set to complement the downstream dataset-specific information through fine-tuning.
We set up a randomly initialized register $\mathbf{E}\in\mathbb{R}^{H\times D_{\mathrm{r}}}$ with $H$ cluster center vectors $\mathbf{e}_{i}\in\mathbb{R}^{D_{\mathrm{r}}},i\in\{1,2,…,H\}$. Each of input time series $\mathbf{x}\in\mathbb{R}^{L}$ is projected into a data-dependent embedding $\mathbf{x}_{\mathrm{e}}\in\mathbb{R}^{D_{\mathrm{r}}}$ through a linear layer.
Pre-training stage. As shown in Figure 2(b), we use the register to cluster these data-dependent embeddings, which generate domain-specific information, and store them in pre-training. Specifically, We find a cluster center vector $\mathbf{e}_{\delta}$ from the register $\mathbf{E}$ where we use $\delta$ to denote the cluster that the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ belongs to.
$\mathcal{L}_{\text{register}}=\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{\delta}\|^{2}_{2},\ \ \delta=\underset{j=1:H}{\arg\operatorname{min}}\left\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{j}\right\|_{2}.$ (6)
To update the cluster center vectors in the register $\mathbf{E}$ that represents the domain information of the pre-trained datasets, we set the loss function shown in Equation 6 that minimizes the distance between the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$. To solve the problem that the gradient of the $\arg\operatorname{min}$ function cannot be backpropagated, we use the stop gradient operation to pass the gradient of $\mathbf{e}_{\delta}$ directly to $\mathbf{x}_{\mathrm{e}}$.
In this way, the vectors in the register $\mathbf{E}$ cluster the embeddings of different data and learn the domain-specific centers for pre-trained datasets, which can represent domain-specific information. As a vector in the register $\mathbf{E}$, $\mathbf{e}_{\delta}$ represents the domain-specific information for input $\mathbf{x}$. $\mathbf{e}_{\delta}$ is invariant under small perturbations in $\mathbf{x}_{\mathrm{e}}$ that represents $\mathbf{x}$, which promotes better representation of domain-specific information and robustness of the vectors in the register. This also avoids their over-reliance on detailed information about specific datasets.
@@
Fine-tuning stage. As shown in Figure 2(c), after obtaining a register $\mathbf{E}$ that contains domain-specific information through pre-training, we freeze the register parameters to adaptively use this domain-specific information in the downstream tasks.
Since the target domain may not strictly fit one of the upstream domains, we propose a novel embedding learning of the downstream data by employing a Top-K strategy that selects $k$ similar vectors in the register. As shown in Equation 7, the embedding of input time series $\mathbf{x}_{\mathrm{e}}$ picks the $k$ nearest vectors in the register $\mathbf{E}$, and uses their average as $\bar{\mathbf{e}}_{k}$ to represent the domain-specific information from the pre-train stage. $\bar{\mathbf{e}}_{k}$ is also patched into $\mathbf{X}_{\mathrm{d}}\in\mathbb{R}^{N_{\text{r}}\times D}$ and is used as the domain specific register tokens.
@@
Since the downstream data has its own specific information at the dataset level in addition to the domain level, this may not be fully represented by the domain information obtained from the pre-trained dataset alone. Therefore, we innovatively set a learnable matrix $\mathbf{A}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ to adjust $\mathbf{X}_{\mathrm{d}}$ to complement the specific information of downstream data. Since the pre-trained model has a very low intrinsic dimension (Aghajanyan et al., 2020), in order to get better fine-tuning results, $\mathbf{A}$ is set as a low-rank matrix:
@@ 3.4 Training
To improve the model’s prediction ability in zero-shot and few-shot setting, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency masking to learn unified features that are more applicable to the downstream prediction task. We normalize time series by employing the REVIN (Kim et al., 2021) that is commonly used by the state-of-the-art time series models, which first normalizes each input sample and subsequently applies inverse normalization to recover the model’s output.
Prediction task. The input time series $\mathbf{x}\in\mathbb{R}^{L}$ is sliced into $P$ non-overlapping patches and then mapped to $\mathbf{X}_{\mathrm{p}}\in\mathbb{R}^{P\times D}$. Based on common forecasted needs (Qiu et al., 2024), we set up four prediction heads mapping to prediction lengths of $\{96,192,336,720\}$ to accomplish the prediction task. Patch tokens $\mathbf{X}_{\mathrm{p}}$ are concatenated with the register tokens $\mathbf{X}_{\mathrm{u}}$ and then successively fed into the Transformer encoder to yield the representation $\mathbf{S}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$:
$\mathbf{S}=\mathrm{Encoder(Concatenate(\mathcal{X}_{p},\mathcal{X}_{u}))}.$ (10)
We feed the representation $\mathbf{S}$ into the prediction decoder and prediction heads to obtain four prediction results $\mathbf{\hat{Y}}_{F}$, where $F\in\{96,192,336,720\}$. With the ground truth $\mathbf{Y}_{F}$, the prediction loss $\mathcal{L}_{\text{prediction}}$ is shown in Equation 11.
@@
Pre-training. The reconstruction task learns generalized features through the Transformer encoder and reconstruction decoder. To utilize these features for the prediction task, the parameters of the reconstruction decoder are copied to the prediction decoder during forward propagation. To avoid prediction training affecting the generalization performance of the model, the gradients of the prediction heads are skipped at back-propagation. The overall loss of ROSE in pre-training stage is shown in Equation 12.
@@ 4 Experiments
Pre-training datasets. The datasets are crucial for pre-training a general time series forecasting model. In light of this, we gather a considerable amount of publicly available datasets from various domains, including energy, nature, health, transport, web, and economics, etc. The details of these datasets are shown in the Appendix A.1.1. To enhance data utilization, we downsample fine-grained datasets to coarser granularity, resulting in approximately 887 million time points.
@@
Evaluation datasets. To conduct comprehensive and fair comparisons for different models, we conducted experiments on seven well-known forecasting benchmarks as the target datasets, including Weather, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
Baselines. We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023), and two LLM-based models: GPT4TS (Zhou et al., 2024) and $\textbf{S}^{2}$-LLM (Pan et al., 2024). In addition, we select five foundation models for comparison in zero-shot setting, including Timer (Liu et al., 2024), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2023b), and Moment (Goswami et al., 2024).
@@
Setup. Consistent with previous works, we adopted Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. For fair comparison, all methods fix the look-back window $L$ = 512 and predict the future values with lengths $F=\{96,192,336,720\}$. More implementation details are presented in the Appendix A.1.3.
4.1 In distribution forecasting
Setting. In full-shot setting, we fine-tune pre-trained ROSE and the baselines with the full downstream data. In few-shot setting, we fine-tune all models with only 10% train data.
@@ 4.1 In distribution forecasting
Full-shot results. As shown in Table 1, we also present the results of the ROSE in 10% few-shot setting. Key observations are summarized as follows. First, as a general forecasting model, ROSE achieves superior performance compared to the six state-of-the-art baselines with full-data training, achieving an average MSE reduction of 15%, which shows that our decomposed frequency learning and register help to learn generalized representations from large-scale datasets and adaptively transfer the multi-domain information to specific downstream scenarios. Second, we observe that ROSE in 10% few-shot setting shockingly improves a large margin as MSE reduction in average exceeding 12% over the baselines trained with full data. This observation validates the transferability of ROSE pre-trained with large multi-source data.
Models ROSE ROSE (10%) ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$-LLM
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.391 0.414 0.397 0.419 0.439 0.448 0.413 0.434 0.582 0.533 0.416 0.436 0.427 0.426 0.406 0.427
ETTh2 0.331 0.374 0.335 0.380 0.374 0.406 0.331 0.381 0.409 0.438 0.508 0.485 0.354 0.394 0.347 0.391
ETTm1 0.341 0.367 0.349 0.372 0.362 0.391 0.353 0.382 0.490 0.464 0.356 0.378 0.352 0.383 0.343 0.379
ETTm2 0.246 0.305 0.250 0.308 0.269 0.329 0.256 0.317 0.317 0.358 0.259 0.325 0.266 0.326 0.257 0.319
Weather 0.217 0.251 0.224 0.252 0.233 0.271 0.226 0.264 0.329 0.336 0.239 0.289 0.237 0.270 0.222 0.259
Electricity 0.155 0.248 0.164 0.253 0.164 0.261 0.159 0.253 0.195 0.296 0.166 0.267 0.167 0.263 0.161 0.257
Traffic 0.390 0.264 0.418 0.278 0.397 0.282 0.391 0.264 0.623 0.333 0.433 0.305 0.414 0.294 0.405 0.286
Few-shot results. The results under the 10% few-shot setting are presented in Table 13 in Appendix A.10.2. ROSE outperforms advanced models when training data is scarce in the target domain.
@@
Figure 4 shows the performance of pre-trained ROSE and ROSE trained from scratch on ETTh1 and ETTm2 with different fine-tuning data percentages, noting the best baselines in full-shot setting. The pre-trained ROSE shows stable, superior performance even with limited fine-tuning samples. Specifically, the pre-trained ROSE exceeds SOTA performance with only 1% train data for ETTh1 and 2% for ETTm2. Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in predictive performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through pre-training.
4.2 Zero-shot forecasting
Setting. In this section, to ensure a fair comparison, we conduct zero-shot predictions for each foundational model on downstream datasets not included in their pre-training data. It is worth noting that, unlike a few foundation models (Woo et al., 2024) that require much longer inputs to achieve better predictive performance, we fix the input length of all baselines to 512 without considering longer input lengths, as many real-world scenarios could offer very limited samples.
Results. As shown in Table 2, ROSE significantly outperforms across the majority of datasets, achieving an average reduction of 15% in Mean Squared Error (MSE). In comparison to Timer and Moirai, ROSE achieves average MSE reductions of 9% and 6%, respectively, and demonstrates a remarkable 43% relative improvement over Moment. Notably, ROSE stands out not only for its superior performance but also for its exceptional lightweight and efficient design, which sets it apart from other foundational models. Detailed analysis of these aspects will be presented in Section 4.3.
Table 2: The results for ROSE and other foundation models in zero-shot setting. The average results of all predicted lengths are listed here. We use ’-’ to indicate that the dataset has been involved in the model’s pre-training, and thus not used for testing.
Models ROSE Timer MOIRAI Chronos TimesFM Moment
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.401 0.425 0.451 0.463 0.475 0.443 0.560 0.452 0.489 0.444 0.708 0.580
ETTh2 0.346 0.394 0.366 0.408 0.379 0.396 0.392 0.397 0.396 0.405 0.392 0.430
ETTm1 0.525 0.471 0.544 0.476 0.714 0.507 0.636 0.495 0.434 0.419 0.697 0.555
ETTm2 0.299 0.352 0.360 0.386 0.343 0.356 0.313 0.363 0.320 0.353 0.319 0.360
Weather 0.265 0.305 0.292 0.312 0.267 0.300 0.288 0.310 - - 0.291 0.323
Electricity 0.234 0.320 0.297 0.375 0.241 0.328 0.245 0.312 - - 0.861 0.766
Traffic 0.588 0.412 0.613 0.407 - - 0.371 0.370 - - 1.411 0.804
4.3 Model Analysis
Efficiency analysis. To exhibit the performance and efficiency advantages of ROSE, we compare its parameter count to other foundation models and evaluate their performance and testing time averaged on ETTh1 and ETTh2 datasets in zero-setting. Similarly, for each specific model, we evaluate its parameter count as well as its performance in full-shot setting and training-to-testing time averaged on the same datasets. Specific implementation details and results can be found in Appendix A.2.
Figure 5: Model performance, number of parameter and efficiency comparison.
As shown in Figure 5, ROSE is a lightweight general model with 7.4M parameters and short inference time, which are only about one-tenth of the second fastest/smallest foundation model (Timer). Importantly, ROSE uses the least number of parameters among foundation models, with its parameter count approaching that of specific models, while exhibiting superior zero-shot performance. This is attributed to our proposed decomposed frequency learning that enhances the model’s comprehension of time series. Concurrently, the TS-Register achieves the adaptive transfer thus efficiently adapting to downstream tasks without the need of scaling up to achieve strong generalizability. Compared to foundation models with large scale, ROSE may better meet the need for general models in real scenarios that require high computational and parameter efficiency as well as high prediction accuracy with scarce downstream data.
Figure 6: Visualization of TS-Register. The calculation of cosine similarity is in the Appendix A.3.
Scalability and Sensitivity. The scalability analysis of ROSE’s model size and pre-training data size are presented in Appendix A.4. The sensitivity analyses for the upper bound $a$ of the thresholds, the number of masked series $K_{\mathrm{f}}$, the number of register tokens $N_{\text{r}}$, the size of register $H$ and number of selections $k$ in Top-K strategy are presented in Appendix A.5.
@@ 4.4 Ablation studies
Masking method. To further validate the effectiveness of decomposed frequency learning, we replace the multi-frequency masking with different masking methods, including two mainstream time-domain methods: patch masking (Nie et al., 2022) and multi-patch masking (Dong et al., 2024), as well as random frequency masking (Chen et al., 2023b). The results in Table 4 show that random frequency masking and patch masking led to negative transfer on ETTm1 and ETTm2, likely due to significant disruption of the original time series, causing overfitting. In contrast, multi-patch masking and multi-frequency masking resulted in positive transfer across all datasets by preventing excessive disruption. Multi-frequency masking achieved better results, demonstrating its ability to help the model understand temporal patterns from a multi-frequency perspective. We also compare with some other pre-training tasks in Table 14 in Appendix A.10.3.
Table 3: Ablations on key components of model architecture, including TS-register, prediction task and reconstruction task. The average results of all predicted lengths are listed here.